# On the Value of Tokeniser Pretraining in Physics Foundation Models

### 05 Mar 2026 v1

@@ Abstract
We investigate the impact of tokeniser pretraining on the accuracy and efficiency of physics emulation. Modern high-resolution simulations produce vast volumes of data spanning diverse physical regimes and scales. Training foundation models to learn the dynamics underlying such data enables the modelling of complex multiphysics phenomena, especially in data-limited settings. The emerging class of physics foundation models typically aims to learn two tasks jointly: (i) extracting compact representations of high-resolution spatiotemporal data, and (ii) capturing governing physical dynamics. However, learning both tasks from scratch simultaneously can impede the effectiveness of either process. We demonstrate that pretraining the tokeniser with an autoencoding objective prior to training the dynamics model enhances computational efficiency for downstream tasks. Notably, the magnitude of this benefit depends on domain alignment: pretraining on the same physical system as the downstream task yields the largest improvements, while pretraining on other systems provides moderate gains. In-domain pretraining reduces VRMSE by 64% after 10,500 training steps compared to training from scratch. To our knowledge, this is the first systematic investigation of tokeniser pretraining for physics foundation models. We further introduce flexible spatiotemporal compression operations that extend causal convolutions to support runtime-adjustable compression ratios, enabling efficient adaptation to diverse downstream tasks. Our findings provide practical guidance for training efficient physics emulators and highlight the importance of strategic pretraining data selection.
@@ 1 Introduction
Modelling physical systems stands to benefit significantly from large-scale pretraining. Foundation models trained across diverse physical systems can capture shared patterns that span different physical regimes and scales; patterns that are often beyond the scope of any single simulation. However, the high spatial and temporal resolution of scientific data makes it computationally prohibitive to train large-scale transformer-based models on the simulations’ pixel space. This underscores the need efficient approaches to processing such data.
In state-of-the-art generative systems operating on high-resolution data, the modelling process is often decomposed into two stages: (i) Tokenisation: extracting compact, informative representations of temporally and spatially high-frequency features using convolutional layers [^1] Here, ”tokenisation” refers to learning representations (i.e., tokens) from data that will be used to train another network; it does not imply that the tokens are quantised or discrete., and (ii) Prediction: feeding these representations into a transformer-based network to model global dependencies and make predictions (NVIDIA et al., 2025; Yu et al., 2024; Tschannen et al., 2024; Yu et al., 2023; Chang et al., 2022).
This separation is motivated by the hypothesis that low-level (i.e., high-frequency) structure is best captured by locally connected architectures - namely, convolutions. At higher semantic levels (i.e., lower frequencies), global interactions become more important, making transformer architectures more suitable (Esser et al., 2021). This hypothesis is supported by observations that transformers tend to learn convolution-like structures at small scales (Dosovitskiy et al., 2021), motivating the explicit use of convolutional layers for encoding high-frequency information. Convolutions are also significantly more efficient to train than transformers due to their local connectivity, parameter sharing, and linear scaling with input size. A growing body of work supports training the tokeniser and transformer in two separate stages rather than jointly from scratch, and this staged approach has become standard in high-resolution image and video generation (NVIDIA et al., 2025; Yu et al., 2024; Rombach et al., 2022).
@@
Pretraining the tokeniser separately offers several potential advantages. First, it decouples representation learning from dynamics modelling, potentially allowing each task to be learned more effectively. Second, a pretrained tokeniser can be reused across multiple downstream applications that require the same latent space, amortising the pretraining cost across different physics problems.
Our contribution. To our knowledge, this is the first systematic investigation of tokeniser pretraining for physics foundation models. We demonstrate that pretraining the tokeniser improves training efficiency for autoregressive prediction of physical dynamics, with benefits dependent on domain alignment between pretraining and downstream tasks. Specifically, in-domain pretraining (same physical system) yields larger improvements than out-of-domain pretraining (different systems). Our focus is on understanding the relative benefits of different pretraining strategies. We prioritise experimental efficiency to enable systematic comparison across pretraining conditions.
Additionally, we introduce flexible spatiotemporal compression operations that extend causal convolutions to support runtime-adjustable compression ratios. This enables the computational cost of downstream tasks to be adjusted dynamically by varying the temporal and spatial coarseness of tokens. Importantly, different physical systems exhibit varying degrees of compressibility - for example, galaxy images compress more readily than turbulent fluid simulations - and flexible compression accommodates this diversity, enabling broader applicability and better adaptation to diverse downstream tasks without retraining.
The rest of this paper is organised as follows. Section 2 describes our methodology, including the datasets, evaluation metrics, model design, and training procedures. Section 3 presents our results on the impact of tokeniser pretraining in both in-domain and out-of-domain settings. Section 4 discusses the implications of our findings, practical considerations for training physics foundation models, and directions for future work. Section 5 provides concluding remarks. Appendix A provides detailed dataset descriptions, Appendix B describes the tokeniser architecture and flexible compression operations, and Appendix C contains complete training hyperparameters.
@@ 2 Methods
This section is structured as follows. First, we describe the datasets used in our experiments (Subsection 2.1). Next, we outline the experimental design and evaluation metrics used to quantify the impact of tokeniser pretraining (Subsection 2.2). We then present the training objectives for both the downstream model and tokeniser (Subsection 2.3), followed by introducing the model architectures (Subsection 2.4). Finally, we detail the parameter counts and training procedures (Subsection 2.5).
@@ 2.2 Experiments and Evaluation Criteria
The primary objective of this study is to investigate whether pretraining the tokeniser provides measurable benefits for downstream foundation model training. To this end, we conduct controlled experiments comparing autoregressive rollout models trained with pretrained tokenisers against those with randomly initialised tokenisers (Figure 1).
Figure 1: Experiments setup.
For the pretrained models, we consider two settings: in-domain, where the tokeniser and rollout model train on the same dataset (Euler), and out-of-domain, where the tokeniser pretrains on a mixture of other datasets (Rayleigh-Bénard, active matter, and shear flow) before the rollout trains on Euler. The latter tests whether pretraining benefits generalise across physical systems.
We evaluate the impact of tokeniser pretraining on training cost and model performance, asking: What is the gain, in terms of compute or quality, when using a pretrained tokeniser? Specifically, does pretraining reduce the computation required to reach a target performance level, or improve model quality under a fixed computational budget? We focus our analysis on early downstream training stages, where practitioners with limited computational resources would most benefit from training efficiency improvements.
@@
$\mathrm{VRMSE}=\frac{\left\lVert\mathbf{x}-\hat{{\mathbf{x}}}\right\rVert_{2}}{\sigma_{{\mathbf{x}}}},$ (1)
where $\sigma_{\mathbf{x}}$ is the standard deviation of the target field $\mathbf{x}$} (computed over all pixels). This measures reconstruction error relative to the variability in the target. We compute VRMSE per timeframe and per field, then average across both to obtain aggregate metrics.
@@
This power spectrum quantifies the contribution of each wavenumber band to the total pixel-space MSE via Plancherel’s theorem: the total MSE equals $\frac{1}{N^{2}}\sum_{k}|B_{k}|\,P_{{\mathbf{x}}-\hat{{\mathbf{x}}}}(k)$, where $N^{2}$ is the number of spatial pixels. To obtain a scale-invariant quality metric, we normalise by the true field’s power spectrum at each scale:
@@
where $P_{{\mathbf{x}}}(k)$ is the power spectrum of the true field, computed analogously to $P_{{\mathbf{x}}-\hat{{\mathbf{x}}}}(k)$. This normalised error power spectrum (NEPS) measures the ratio of error power to signal power at each scale. Since each frequency band uses its own normalisation, NEPS values are not additive across scales but provide a relative quality metric independent of the natural energy distribution. A value of NEPS(k)=0.1(k)=0.1 indicates the error power is 10% of the signal power at scale $k$, while NEPS(k)=1.0(k)=1.0 means the error power equals the signal power at that scale. We partition wavenumbers into three frequency ranges (low, mid, high) and report NEPS for each range. We compute NEPS per timeframe and per field, obtaining aggregate metrics by averaging across timeframes and/or fields.
@@ 2.3 Training Objective
$\ell_{\text{Rollout}}=\left\lVert{\mathbf{x}}_{t}-\hat{{\mathbf{x}}}_{t}\right\rVert_{1}$ (4)
@@
$\ell_{\text{Tokeniser}}=\left\lVert{\mathbf{x}}_{0:t-1}-\hat{{\mathbf{x}}}_{0:t-1}\right\rVert_{2}^{2}$ (5)
@@ 2.4 Architecture
During downstream rollout training, the processor and the projection layer are always trainable. For the tokeniser, we consider two parameter training strategies. In the fully trainable setting, all tokeniser parameters are updated jointly with the processor. In the mostly frozen setting, the core of the tokeniser is frozen and only the layers directly interfacing with pixel space and latent space are trainable. Specifically, this includes the encoder head, the decoder head, and the bottleneck layers, while all intermediate blocks remain frozen.
@@ 2.5 Parameter Counts and Training Recipe
The tokeniser’s encoder progressively increases channel dimensions from the number of physical fields to 16, 32, and 64 channels, then compresses to a bottleneck of 18 latent channels. The decoder mirrors this structure. The processor consists of six transformer blocks with embedding dimension 1088. Table 2 summarises the parameter counts for each component. Table 3 shows the number of trainable parameters under each parameter training strategy. In the fully trainable setting, all tokeniser parameters are updated during rollout training. In the mostly frozen setting, only 2% of the tokeniser remain trainable.
@@ 3.2 Downstream Performance
Table 5 reports model performance on next-frame prediction after 10,500 training steps (168,000 examples across eight GPUs). We focus on this early training regime to assess the computational efficiency benefits of pretraining under limited budgets. All metrics are evaluated on the validation set by computing predictions for frame $t$ given context frames 0 through t1t-1, then comparing the predicted frame $\hat{x}_{t}$ to the ground truth $x_{t}$. To reduce sensitivity to the particular choice of validation examples, we randomly sample 10 subsets of 4096 examples from the validation set, compute metrics on each subset, and report the mean across these 10 subsets. Figure 2 shows how these validation metrics evolve over a period of 29,400 steps for a representative run.
Table 5: Downstream task performance on the validation set after 10,500 training steps under different pretraining and freezing strategies. Metrics are averaged over 10 random subsets of 4096 validation examples. The
@@
Figure 2: Next-frame prediction learning curves on the validation set over 29,400 training steps for different pretraining and freezing configurations. Each panel shows a different evaluation metric: VRMSE (leftmost), and spectral error at low, mid, and high frequency ranges.
Tokeniser pretraining provides substantial improvements when training with limited computational budgets, with gains depending critically on domain alignment. In-domain pretraining reduces VRMSE by 64% compared to training from scratch (Table 5: 0.158 vs 0.439 for fully trainable; 0.162 vs 0.439 when mostly frozen). Out-of-domain pretraining provides more modest gains of 19% when the tokeniser remains trainable, though freezing it in this setting degrades performance.
Figure 2 reveals how pretraining affects learning dynamics across spatial scales. In-domain pretrained models rapidly achieve the lowest VRMSE, with both fully trainable and mostly frozen variants reaching low error within 10,000 steps and maintaining this advantage throughout training. At low frequencies, in-domain models show dramatic improvement of over two orders of magnitude, while out-of-domain trainable and no-pretraining models eventually reach similar error levels, and the out-of-domain frozen model remains at higher error. At mid frequencies, all pretrained models initially outperform training from scratch, with the advantage persisting longest for in-domain pretraining: the out-of-domain trainable model maintains lower error through approximately 15,000 steps before converging to similar performance as no pretraining, while the out-of-domain frozen model plateaus at intermediate error levels. At high frequencies, all models show poor quality (NEPS $\approx$ 1.0), but their learning dynamics differ: in-domain pretraining shows improving performance, whereas models without pretraining or with out-of-domain pretraining show degrading performance over training.
@@ 3.3 Freezing Strategies
For the in-domain case, the fully trainable and mostly frozen variants achieve nearly identical performance on the first prediction step, raising the question of whether one approach offers any advantages. To investigate this, we examine later rollout steps, where errors can accumulate, and extend training to 210,000 steps to assess whether differences emerge with longer training. Specifically, we evaluate autoregressive rollout quality over 18 prediction steps. [^2] Starting from 9 frames of ground truth context ($t=0$ to $t=8$), the model autoregressively predicts the next frame ($t=9$). Each prediction is then fed back as input for the subsequent prediction. We continue this process until frame $t=26$ (18 autoregressive steps). Figure 3 shows VRMSE on the validation set across early (steps 1-2), medium (steps 3-6), and long (steps 7-18) rollout horizons for a representative run, with each metric averaged over 1024 validation examples.
Figure 3: Autoregressive rollout learning curves over 210,000 training steps at different prediction horizons (single run). Each panel shows VRMSE for a different rollout window. The shaded region indicates the training range shown in Figure 2.

### 11 Mar 2026 v2

@@ Abstract
We investigate the impact of tokeniser pretraining on the accuracy and efficiency of physics emulation. Modern high-resolution simulations produce vast volumes of data spanning diverse physical regimes and scales. Training foundation models to learn the dynamics underlying such data enables the modelling of complex multiphysics phenomena, especially in data-limited settings. The emerging class of physics foundation models typically aims to learn two tasks jointly: (i) extracting compact representations of high-resolution spatiotemporal data, and (ii) capturing governing physical dynamics. However, learning both tasks from scratch simultaneously can impede the effectiveness of either process. We show that pretraining the tokeniser with an autoencoding objective prior to training the dynamics model enhances computational efficiency for physics emulation. Notably, the magnitude of this benefit depends on domain alignment: pretraining on the same physical system as the emulation task yields the largest improvements, while pretraining on other systems provides moderate gains. In-domain pretraining reduces VRMSE by 64% after 10,500 training steps compared to training from scratch. To our knowledge, this is the first systematic investigation of tokeniser pretraining for physics foundation models. We further introduce flexible spatiotemporal compression operations that extend causal convolutions to support runtime-adjustable compression ratios, enabling efficient adaptation to diverse downstream tasks. Our findings provide practical guidance for training efficient physics emulators and highlight the importance of strategic pretraining data selection.
@@ 1 Introduction
Modelling physical systems stands to benefit significantly from large-scale pretraining. Foundation models trained across diverse physical systems can capture shared patterns that span different physical regimes and scales; patterns that are often beyond the scope of any single simulation. However, the high spatial and temporal resolution of scientific data makes it computationally prohibitive to train large-scale transformer-based foundation models directly in pixel space. This underscores the need for efficient approaches to processing such data.
In state-of-the-art generative systems operating on high-resolution data, the modelling process is often decomposed into two stages: (i) Tokenisation: extracting compact, informative representations of temporally and spatially high-frequency features using convolutional layers, [^1] Here, “tokenisation” refers to learning representations (i.e., tokens) from data that will be used to train another network; it does not imply that the tokens are quantised or discrete. and (ii) Prediction: feeding these representations into a transformer-based network to model global dependencies and make predictions (NVIDIA et al., 2025; Yu et al., 2024; Tschannen et al., 2024; Yu et al., 2023; Chang et al., 2022).
This separation is motivated by the hypothesis that low-level (i.e., high-frequency) structure is best captured by locally connected architectures – namely, convolutions. At higher semantic levels (i.e., lower frequencies), global interactions become more important, making transformer architectures more suitable (Esser et al., 2021). This hypothesis is supported by observations that transformers tend to learn convolution-like structures at small scales (Dosovitskiy et al., 2021), motivating the explicit use of convolutional layers for encoding high-frequency information. Convolutions are also significantly more efficient to train than transformers due to their local connectivity, parameter sharing, and linear scaling with input size. A growing body of work supports training the tokeniser and transformer in two separate stages rather than jointly from scratch, and this staged approach has become standard in high-resolution image and video generation (NVIDIA et al., 2025; Yu et al., 2024; Rombach et al., 2022).
@@
Pretraining the tokeniser offers several potential advantages. First, it decouples representation learning from dynamics modelling, allowing each task to be learned more effectively. Second, a pretrained tokeniser can be reused across multiple downstream applications that share the same latent space, amortising the pretraining cost across different physics problems.
Our contribution. To our knowledge, this is the first systematic study of tokeniser pretraining for physics foundation models. We demonstrate that pretraining the tokeniser improves training efficiency for autoregressive prediction of physical dynamics, with benefits dependent on domain alignment between pretraining data and autoregressive training data. Specifically, in-domain pretraining (same physical system) yields larger improvements than out-of-domain pretraining (different systems). Our focus is on understanding the relative benefits of different pretraining strategies. We therefore design our experiments to facilitate systematic comparisons across pretraining conditions.
Additionally, we introduce flexible spatiotemporal compression operations that extend causal convolutions to support runtime-adjustable compression ratios. This enables the computational cost of downstream tasks to be adjusted dynamically by varying the temporal and spatial coarseness of tokens. Importantly, different physical systems exhibit varying degrees of compressibility – for example, galaxy images compress more readily than turbulent fluid simulations – and flexible compression accommodates this diversity, enabling broader applicability and better adaptation to diverse downstream tasks without retraining.
The rest of this paper is organised as follows: Section 2 describes our methodology, including the datasets, evaluation metrics, model design, and training procedures. Section 3 presents our results on the impact of tokeniser pretraining in both in-domain and out-of-domain settings. Section 4 discusses the implications of our findings, practical considerations for training physics foundation models, and directions for future work. Section 5 provides concluding remarks. Appendix A provides detailed dataset descriptions, Appendix B describes the tokeniser architecture and flexible compression operations, and Appendix C contains complete training hyperparameters.
@@ 2 Methods
This section is structured as follows:  First, we describe the datasets used in our experiments (Subsection 2.1). Next, we outline the experimental design and evaluation metrics (Subsection 2.2). We then present the training objectives for both the dynamics model and tokeniser (Subsection 2.3), followed by introducing the model architectures (Subsection 2.4). Finally, we detail the parameter counts and training procedures (Subsection 2.5).
@@ 2.2 Experiments and Evaluation Criteria
The primary objective of this study is to investigate whether pretraining the tokeniser provides measurable benefits for foundation model training. To this end, we conduct controlled experiments comparing autoregressive rollout models trained with pretrained tokenisers against those with randomly initialised tokenisers (Figure 1).
Figure 1: Overview of the experimental setup.
For the pretrained models, we consider two settings: in-domain, where both the tokeniser and rollout model are trained on the same dataset (Euler), and out-of-domain, where the tokeniser is pretrained on a mixture of other datasets (Rayleigh-Bénard, active matter, and shear flow) before the rollout model is trained on Euler. The latter tests whether the benefits of pretraining generalise across physical systems. Aside from the tokeniser, the rest of the rollout model is randomly initialised in all settings.
We evaluate the impact of tokeniser pretraining on training cost and model performance, asking: What is the gain, in terms of compute or quality, when using a pretrained tokeniser? Specifically, does pretraining reduce the computation required to reach a target performance level, or improve model quality under a fixed computational budget? We focus our analysis on early stages of downstream training, where practitioners with limited computational resources would most benefit from training efficiency improvements. Training cost is quantified in wall-clock time. We evaluate model performance using two complementary metrics:
@@
$\mathrm{VRMSE}=\frac{\sqrt{\mathbb{E}_{p}\!\bigl[(x_{p}-\hat{x}_{p})^{2}\bigr]}}{\sigma_{{\mathbf{x}}}},$ (1)
where $\mathbf{x}$} and $\hat{{\mathbf{x}}}$ denote the target and predicted fields, $p$ indexes pixels, and $\sigma_{\mathbf{x}}$ is the standard deviation of the target field computed over all pixels. This measures reconstruction error relative to the variability in the target. We compute VRMSE per timeframe and per field, then average across both to obtain aggregate metrics.
@@
This power spectrum quantifies the contribution of each wavenumber band to the total pixel-space MSE via Plancherel’s theorem: the total MSE equals $\frac{1}{N^{2}}\sum_{k}|B_{k}|\,P_{{\mathbf{x}}-\hat{{\mathbf{x}}}}(k)$, where $N^{2}$ is the number of spatial pixels. To obtain a scale-invariant metric, we normalise by the true field’s power spectrum at each scale:
@@
where $P_{{\mathbf{x}}}(k)$ is the power spectrum of the true field, computed analogously to $P_{{\mathbf{x}}-\hat{{\mathbf{x}}}}(k)$. This normalised error power spectrum (NEPS) measures the ratio of error power to signal power at each scale. Since each frequency band uses its own normalisation, NEPS values are not additive across scales but provide a relative quality metric independent of the natural energy distribution. A value of NEPS(k)=0.1(k)=0.1 indicates the error power is 10% of the signal power at scale $k$, while NEPS(k)=1.0(k)=1.0 means the error power equals the signal power at that scale. We partition wavenumbers into three frequency ranges (low, mid, high) and report NEPS for each range. We compute NEPS per timeframe and per field, obtaining aggregate metrics by averaging across timeframes and fields.
@@ 2.3 Training Objective
$\ell_{\text{Rollout}}=\left\lVert{\mathbf{x}}_{t}-\hat{{\mathbf{x}}}_{t}\right\rVert_{1}.$ (4)
@@
$\ell_{\text{Tokeniser}}=\left\lVert{\mathbf{x}}_{0:t-1}-\hat{{\mathbf{x}}}_{0:t-1}\right\rVert_{2}^{2}.$ (5)
@@ 2.4 Architecture
During rollout training, the processor and the projection layer are always trainable. For the tokeniser, we consider two parameter training strategies. In the fully trainable setting, all tokeniser parameters are updated jointly with the processor. In the mostly frozen setting, the core of the tokeniser is frozen and only the layers directly interfacing with pixel space and latent space are trainable. Specifically, this includes the encoder head, the decoder head, and the bottleneck layers, while all intermediate blocks remain frozen.
@@ 2.5 Parameter Counts and Training Recipe
The tokeniser’s encoder progressively increases channel dimensions from the number of physical fields to 16, 32, and 64 channels, then compresses to a bottleneck of 18 latent channels. The decoder mirrors this structure. The processor consists of six transformer blocks with embedding dimension 1088. Table 2 summarises the parameter counts for each component. Table 3 shows the number of trainable parameters under each parameter training strategy. In the fully trainable setting, all tokeniser parameters are updated during rollout training. In the mostly frozen setting, only 2% of them remain trainable.
@@ 3.2 Downstream Performance
Table 5 reports the performance of the rollout model on next-frame prediction after 10,500 training steps (168,000 examples across eight GPUs). We focus on this early training regime to assess the computational efficiency benefits of pretraining under limited budgets. All metrics are evaluated on the validation set by computing predictions for frame $t$ given context frames 0 through t1t-1, then comparing the predicted frame $\hat{{\mathbf{x}}}_{t}$ to the ground truth ${\mathbf{x}}_{t}$. To reduce sensitivity to the particular choice of validation examples, we randomly sample 10 subsets of 4096 examples from the validation set, compute metrics on each subset, and report the mean across these 10 subsets. Figure 2 shows how these validation metrics evolve over a period of 29,400 steps for a representative run.
Table 5: Next-frame prediction performance on the validation set after 10,500 training steps for different pretraining and freezing configurations. Metrics are averaged over 10 random subsets of 4096 validation examples. The
@@
Figure 2: Next-frame prediction learning curves on the validation set, shown over 29,400 training steps for different pretraining and freezing configurations. Each panel shows a different evaluation metric: VRMSE (leftmost), and spectral error at low-, mid-, and high-frequency ranges.
Tokeniser pretraining provides substantial improvements under limited computational budgets, with gains depending on domain alignment. In-domain pretraining reduces VRMSE by 64% compared to training from scratch (Table 5: 0.158 vs. 0.439 for fully trainable; 0.162 vs. 0.439 when mostly frozen). Out-of-domain pretraining provides more modest gains of 19% when the tokeniser remains trainable, though freezing it in this setting degrades performance.
Figure 2 reveals how pretraining affects learning dynamics across spatial scales. In-domain pretrained models rapidly achieve the lowest VRMSE: both fully trainable and mostly frozen variants reach low error within 10,000 steps and maintain this advantage throughout training. At low frequencies, in-domain models improve by more than two orders of magnitude early in training. The out-of-domain trainable and no-pretraining models eventually reach similar error levels to each other, while the out-of-domain frozen model remains at higher error. At mid frequencies, all pretrained models initially outperform training from scratch. In-domain pretraining achieves the lowest error and maintains this advantage throughout training. The out-of-domain trainable model converges to the no-pretraining baseline around 20,000 steps, whereas the out-of-domain frozen model begins to underperform the baseline around 15,000 steps. At high frequencies, all models produce poor quality (NEPS $\approx$ 1.0), but their learning dynamics differ: in-domain models show improving performance, whereas models without pretraining or with out-of-domain pretraining degrade over training.
@@ 3.3 Freezing Strategies
For the in-domain case, the fully trainable and mostly frozen variants achieve nearly identical performance on next-frame prediction, raising the question of whether one approach offers any advantages. To investigate this, we examine later rollout steps, where errors can accumulate. We also extend training to 210,000 steps to assess whether differences emerge with longer training. Specifically, we evaluate autoregressive rollout quality over 18 prediction steps. [^2] Starting from 9 frames of ground truth context ($t=0$ to $t=8$), the model autoregressively predicts the next frame ($t=9$). Each prediction is then fed back as input for the subsequent prediction. We continue this process until frame $t=26$ (18 autoregressive steps). Figure 3 shows VRMSE on the validation set across early (step 1 and step 2), medium (steps 3-6), and long (steps 7-18) rollout horizons for a representative run, with each metric averaged over 1024 validation examples.
Figure 3: Autoregressive rollout learning curves over 210,000 training steps. Each panel shows VRMSE at a different rollout horizon. The shaded region indicates the training range in Figure 2.